if a stretch of mountain road covers a horizontal didtance of 0.5 mile and has a veryical drop of 20 feet, find the prtcent grade of the road and the length of thr actual stretch of road.
Answer by addingup(3677)

The % grade is: rise/run = 20/2640 = 0.76%
sqrt(20^2+2640^2) = sqrt(6970000) = 2640.1 feet is the length of the actual stretch of road, or 0.5 miles